# CLIPArTT: Light-weight Adaptation of CLIP to New Domains at Test Time

### v1

@@ 4 Experimental Settings
We evaluate CLIPArTT’s performance across diverse TTA datasets using established methodologies. These datasets simulate challenging domain shifts, offering insights into our approach’s efficacy. Additionally, we explore CLIPArTT’s adaptability on the original dataset through zero-shot test-time adaptation, crucial given the novel nature of the datasets. Our evaluation framework encompasses distinct categories, namely natural images, common corruptions, simulated shift and video shift, providing a comprehensive assessment of the model’s performance across diverse challenges.
In our evaluation of natural images, we utilize CIFAR-10, CIFAR-10.1, and CIFAR-100 datasets, each comprising 10,000 images and featuring 10 and 100 classes, respectively. These datasets represent natural imagery and are novel to the model under scrutiny. Notably, CIFAR-10.1 introduces a natural domain shift from CIFAR-10, thereby enriching our assessment with varied and nuanced data distributions.
Transitioning to our investigation of common corruptions, we turn to the CIFAR-10-C and CIFAR-100-C datasets Hendrycks and Dietterich (2019). These datasets offer a diverse range of 15 distinct corruptions, including elastic transform and impulse noise, among others. Each corruption is characterized by 5 severity levels, yielding a total of 75 unique testing scenarios per dataset. Within each severity level, there are 10,000 images, contributing to a comprehensive evaluation of the model’s robustness against a variety of corruption types and intensities.
Furthermore, we examine the challenging VisDA-C dataset Peng et al. (2018), which represents two domain shifts: the simulated shift and video shift. The former comprises a set of 152,397 images rendered in 3D across 12 different classes. The later includes 72,372 YouTube video frames spanning the same categories. This dataset presents an important challenge, as it bridges the gap of the type of imagery that a model can be applied on.
Test-time adaptation.  For test-time adaptation, model updates are applied to all the Layer Normalization (LN) layers within the visual encoder. We employ the ADAM optimizer with a fixed learning rate set to $10^{-3}$. Throughout our experiments, a consistent batch size of 128 is utilized to maintain uniformity and enable effective comparisons across different scenarios. As in previous TTA works, a smaller learning rate of $10^{-4}$ is preferred to adapt on the 3D renderings split Vargas Hakim et al. (2023), as it represents a more aggressive shift.
Benchmarking.  We compare the performance of CLIPArTT with state-of-the-art methods. Specifically, we utilize adapted versions of TENT Wang et al. (2021) and LAME Boudiaf et al. (2022) tailored to CLIP. The adapted TENT now updates only the affine parameters in Layer Normalization layers through entropy minimization directly on the similarity between image and text features, diverging from its original implementation (i.e., using classifier’s logits). Notably, we employ 10 iterations for TENT adaptation, a choice informed by our findings, as evidenced in Fig 2. For VisDA-C, we also explore using only 1 iteration, as in the original paper. Similarly, we utilize the image-to-text similarity scores as logits for LAME, where the Laplacian regularizer is applied to the similarity between image features. Finally, we extend the evaluation to include TPT Shu et al. (2022), a recent adaptation method for CLIP. TPT heavily relies on image augmentations, thus we reduce batch-size of 32 due to its high memory requirement.
@@
Original 89.80 $\pm$0.05 90.04 $\pm$0.13 90.41 $\pm$0.07
CIFAR 10.1 85.37 $\pm$0.17 86.35 $\pm$0.27 86.07 $\pm$0.21
C-Average 70.79 71.17 70.99
@@
$K$ = 3 $K$ = 4 $K$ = 5
3D 84.31 84.24 87.24
YT 82.80 83.17 83.46
Table 3: Accuracy (%) on the VisDA-C 3D training split (3D) and validation split (YT) with ViT-B/32 on different numbers of selected classes ($K$) to create pseudo-labels.
@@ 5.1 Ablation Studies
Selection of the number of classes for new prompts.  Determining the number of classes in building new prompts presents a critical aspect in our methodology. Table 13 illustrates various scenarios where $K\!=\!3$ emerges as a favorable choice, exhibiting a remarkable accuracy of $100\%$ across multiple instances of CIFAR-10 datasets. This initial observation is further corroborated by the findings presented in Table 14. However, the decision becomes more nuanced when applied to CIFAR-100 datasets. While leveraging the top 3 classes contributes to enhanced accuracy, it does not guarantee optimal performance. Similarly, adopting a strategy akin to CIFAR-10, wherein $30\%$ of the classes are chosen, proves impractical due to resulting lengthy sentences that are impossible to tokenize. Examination of Table 4 reveals that the optimal choice varies depending on the distribution shift. Nonetheless, our analysis shows that $K\!=\!3$ consistently yields superior performance, particularly evident in the average results across CIFAR-100-C. Consequently, we maintain $K\!=\!3$ as the preferred configuration for all subsequent experiments on CIFAR datasets. A similar study was conducted for VisDA-C, in which $K\!=\!5$ (See Table 3) is chosen as the number of classes.
$K$ = 1 $K$ = 3 $K$ = 5 $K$ = 7 $K$ = 10 $K$ = 20
Original 69.00 $\pm$0.22 69.79 $\pm$0.04 69.68 $\pm$0.07 69.56 $\pm$0.02 69.78 $\pm$0.02 69.93 $\pm$0.08
Gaussian Noise 26.05 $\pm$0.11 25.32 $\pm$0.14 24.69 $\pm$0.03 24.60 $\pm$0.03 24.70 $\pm$0.15 23.73 $\pm$0.07
Shot noise 28.88 $\pm$0.11 27.90 $\pm$0.05 27.25 $\pm$0.26 26.75 $\pm$0.07 26.83 $\pm$0.26 25.73 $\pm$0.13
Impulse Noise 24.04 $\pm$0.09 25.62 $\pm$0.09 25.12 $\pm$0.14 25.25 $\pm$0.14 24.95 $\pm$0.23 24.57 $\pm$0.12
Defocus blur 49.03 $\pm$0.15 49.88 $\pm$0.23 49.75 $\pm$0.11 49.74 $\pm$0.25 49.62 $\pm$0.10 49.49 $\pm$0.07
Glass blur 26.77 $\pm$0.14 27.89 $\pm$0.03 27.76 $\pm$0.23 27.28 $\pm$0.07 26.57 $\pm$0.07 25.52 $\pm$0.09
Motion blur 46.50 $\pm$0.09 47.93 $\pm$0.14 47.48 $\pm$0.21 47.57 $\pm$0.10 47.53 $\pm$0.09 47.36 $\pm$0.09
Zoom blur 52.08 $\pm$0.12 52.70 $\pm$0.06 52.22 $\pm$0.10 52.10 $\pm$0.24 52.62 $\pm$0.24 52.82 $\pm$0.09
Snow 49.24 $\pm$0.07 49.72 $\pm$0.01 48.98 $\pm$0.08 48.87 $\pm$0.08 49.13 $\pm$0.08 49.54 $\pm$0.13
Frost 49.91 $\pm$0.07 49.63 $\pm$0.12 48.43 $\pm$0.17 48.11 $\pm$0.06 48.72 $\pm$0.08 49.13 $\pm$0.10
Fog 47.15 $\pm$0.04 48.77 $\pm$0.04 48.95 $\pm$0.18 48.78 $\pm$0.22 49.06 $\pm$0.05 48.74 $\pm$0.36
Brightness 60.01 $\pm$0.08 61.27 $\pm$0.08 60.77 $\pm$0.16 60.89 $\pm$0.19 60.98 $\pm$0.18 61.03 $\pm$0.19
Contrast 46.90 $\pm$0.21 48.55 $\pm$0.24 49.01 $\pm$0.14 49.07 $\pm$0.03 49.27 $\pm$0.09 49.08 $\pm$0.12
Elastic transform 36.32 $\pm$0.10 37.45 $\pm$0.08 37.63 $\pm$0.12 37.31 $\pm$0.09 37.13 $\pm$0.16 36.94 $\pm$0.11
Pixelate 32.52 $\pm$0.17 33.88 $\pm$0.14 34.40 $\pm$0.15 34.38 $\pm$0.16 34.65 $\pm$0.09 34.32 $\pm$0.02
JPEG compression 35.81 $\pm$0.11 36.07 $\pm$0.32 35.77 $\pm$0.01 35.60 $\pm$0.10 35.63 $\pm$0.10 35.29 $\pm$0.14
Average 40.75 41.51 41.21 41.09 41.16 40.89
Table 4: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with Level 5 corruption for different number of $K$ selected classes to create pseudo-labels.
@@
Selection of the number of iterations.  In line with our approach for TENT, we investigate optimal number of iterations. As demonstrated in Table 5, iterating 10 times strikes the most favorable balance, exhibiting superior performance across the overall average on CIFAR-10-C and yielding optimal results for numerous corruption types. Nonetheless, it is imperative to acknowledge that employing a lower number of iterations (e.g., Iter = 5) may yield improved outcomes in scenarios with minimal or no distribution shift, while a higher number of iterations (e.g., Iter = 20) may be more effective in mitigating severe distribution shifts. As a result, 10 iterations are used for all forthcoming experiments.
Iter = 1 Iter = 5 Iter = 10 Iter = 20
Original 89.59 $\pm$0.01 90.54 $\pm$0.09 90.04 $\pm$0.13 88.32 $\pm$0.12
CIFAR 10.1 84.78 $\pm$0.02 86.67 $\pm$0.06 86.35 $\pm$0.27 84.33 $\pm$0.31
Gaussian Noise 39.75 $\pm$0.04 53.79 $\pm$0.08 59.90 $\pm$0.36 59.33 $\pm$0.20
Shot noise 43.80 $\pm$0.04 57.16 $\pm$0.24 62.77 $\pm$0.07 63.08 $\pm$0.43
Impulse Noise 45.19 $\pm$0.07 52.44 $\pm$0.14 56.02 $\pm$0.16 56.73 $\pm$0.01
Defocus blur 72.93 $\pm$0.07 76.36 $\pm$0.10 76.74 $\pm$0.05 75.33 $\pm$0.13
Glass blur 46.61 $\pm$0.06 57.45 $\pm$0.13 61.77 $\pm$0.16 62.01 $\pm$0.27
Motion blur 67.89 $\pm$0.03 74.34 $\pm$0.14 76.01 $\pm$0.19 75.94 $\pm$0.28
Zoom blur 73.24 $\pm$0.05 77.03 $\pm$0.07 77.40 $\pm$0.20 75.42 $\pm$0.13
Snow 73.81 $\pm$0.07 76.51 $\pm$0.08 77.29 $\pm$0.16 76.18 $\pm$0.21
Frost 74.80 $\pm$0.04 78.13 $\pm$0.12 79.20 $\pm$0.08 77.44 $\pm$0.30
Fog 69.81 $\pm$0.06 74.16 $\pm$0.09 75.74 $\pm$0.14 74.66 $\pm$0.02
Brightness 84.16 $\pm$0.01 86.61 $\pm$0.10 86.59 $\pm$0.16 85.14 $\pm$0.42
Contrast 67.75 $\pm$0.03 74.85 $\pm$0.04 77.82 $\pm$0.14 77.75 $\pm$0.11
Elastic transform 63.15 $\pm$0.08 68.53 $\pm$0.19 70.20 $\pm$0.01 68.48 $\pm$0.24
Pixelate 54.20 $\pm$0.02 61.87 $\pm$0.04 66.52 $\pm$0.13 67.13 $\pm$0.15
JPEG compression 57.46 $\pm$0.09 62.00 $\pm$0.13 63.51 $\pm$0.14 63.64 $\pm$0.20
Average 62.30 68.75 71.17 70.55
@@
Image Text Image + Text
Original 90.18 $\pm$0.02 89.05 $\pm$0.14 90.04 $\pm$0.13
CIFAR 10.1 86.25 $\pm$0.37 84.85 $\pm$0.40 86.35 $\pm$0.27
Gaussian Noise 59.05 $\pm$0.30 59.29 $\pm$0.27 59.90 $\pm$0.36
Shot noise 62.11 $\pm$0.18 61.89 $\pm$0.31 62.77 $\pm$0.07
Impulse Noise 55.43 $\pm$0.12 55.48 $\pm$0.10 56.02 $\pm$0.16
Defocus blur 76.88 $\pm$0.09 76.25 $\pm$0.10 76.74 $\pm$0.05
Glass blur 61.56 $\pm$0.03 61.28 $\pm$0.33 61.77 $\pm$0.16
Motion blur 76.32 $\pm$0.15 75.37 $\pm$0.27 76.01 $\pm$0.19
Zoom blur 77.66 $\pm$0.09 76.29 $\pm$0.11 77.40 $\pm$0.20
Snow 77.28 $\pm$0.08 76.03 $\pm$0.16 77.29 $\pm$0.16
Frost 78.80 $\pm$0.18 78.05 $\pm$0.21 79.20 $\pm$0.08
Fog 75.69 $\pm$0.17 73.70 $\pm$0.15 75.74 $\pm$0.14
Brightness 86.62 $\pm$0.20 84.69 $\pm$0.04 86.59 $\pm$0.16
Contrast 77.43 $\pm$0.12 74.52 $\pm$0.02 77.82 $\pm$0.14
Elastic transform 69.63 $\pm$0.02 69.33 $\pm$0.20 70.20 $\pm$0.01
Pixelate 66.33 $\pm$0.16 64.86 $\pm$0.29 66.52 $\pm$0.13
JPEG compression 63.92 $\pm$0.13 63.44 $\pm$0.20 63.51 $\pm$0.14
Average 70.98 70.03 71.17
CLIP BS = 8 BS = 16 BS = 32 BS = 64 BS = 128
C-Average 59.22 61.10 64.72 67.70 69.82 71.17
@@
Performance of the model for different batch-sizes.  TTA methods have historically exhibited limitations when applied with small batch sizes. In our study, we investigate the performance of our model in this regard. As depicted in Table 15, performance improves significantly with increasing batch size. However, beyond a batch size of 8, no further performance gains are observed. This phenomenon can be attributed to the fact that, with smaller batch sizes, our method has the potential to introduce uncertainty by introducing multiple classes, thereby diminishing performance, especially when the model is already confident. In subsequent experiments, we maintain a batch size of 128, consistent with prevailing practices in the state-of-the-art methodologies.
5.2 Experiments
CLIP LAME TENT TPT (BS=32) CLIPArTT
Original 88.74 89.36 $\pm$0.06 91.69 $\pm$0.10 88.06 $\pm$0.06 90.04 $\pm$0.13
CIFAR 10.1 83.25 81.22 $\pm$0.33 87.60 $\pm$0.45 81.80 $\pm$0.27 86.35 $\pm$0.27
Gaussian Noise 35.27 24.71 $\pm$0.11 41.27 $\pm$0.27 33.90 $\pm$0.08 59.90 $\pm$0.36
Shot noise 39.67 27.44 $\pm$0.09 47.20 $\pm$0.23 38.20 $\pm$0.02 62.77 $\pm$0.07
Impulse Noise 42.61 31.38 $\pm$0.15 48.58 $\pm$0.31 37.66 $\pm$0.20 56.02 $\pm$0.16
Defocus blur 69.76 62.45 $\pm$0.44 77.12 $\pm$0.16 67.83 $\pm$0.28 76.74 $\pm$0.05
Glass blur 42.40 29.96 $\pm$0.06 52.65 $\pm$0.30 38.81 $\pm$0.12 61.77 $\pm$0.16
Motion blur 63.97 54.00 $\pm$0.36 71.25 $\pm$0.09 63.39 $\pm$0.13 76.01 $\pm$0.19
Zoom blur 69.83 61.97 $\pm$0.36 76.20 $\pm$0.19 68.95 $\pm$0.16 77.40 $\pm$0.20
Snow 71.78 64.61 $\pm$0.48 78.29 $\pm$0.20 70.16 $\pm$0.10 77.29 $\pm$0.16
Frost 72.86 65.17 $\pm$0.17 79.84 $\pm$0.09 72.39 $\pm$0.22 79.20 $\pm$0.08
Fog 67.04 59.13 $\pm$0.49 77.39 $\pm$0.01 64.31 $\pm$0.28 75.74 $\pm$0.14
Brightness 81.87 80.05 $\pm$0.23 87.78 $\pm$0.03 81.30 $\pm$0.18 86.59 $\pm$0.16
Contrast 64.37 56.91 $\pm$0.37 79.47 $\pm$0.11 62.26 $\pm$0.31 77.82 $\pm$0.14
Elastic transform 60.83 53.89 $\pm$0.20 70.00 $\pm$0.25 56.43 $\pm$0.27 70.20 $\pm$0.01
Pixelate 50.53 39.67 $\pm$0.34 63.74 $\pm$0.18 42.80 $\pm$0.40 66.52 $\pm$0.13
JPEG compression 55.48 47.24 $\pm$0.14 62.64 $\pm$0.14 53.67 $\pm$0.25 63.51 $\pm$0.14
Average 59.22 50.57 67.56 56.80 71.17
Table 8: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with ViT-B/32 as visual encoder.

### v2

@@ 4 Experimental Settings
We evaluate CLIPArTT’s performance across a variety of TTA datasets, covering scenarios such as natural images, common corruptions, simulated images, video, and Domain Generalization benchmark. This comprehensive framework allows for a robust assessment of the model’s adaptability to various challenges, including domain shifts and corruptions. For a detailed description of the datasets, please refer to the supplementary materials.
Test-time adaptation.  For test-time adaptation, model updates are applied to all the Layer Normalization (LN) layers within the visual encoder. We employ the ADAM optimizer with a fixed learning rate set to $10^{-3}$. Throughout our experiments, a consistent batch size of 128 is utilized to maintain uniformity and enable effective comparisons across different scenarios. As in previous TTA works, a smaller learning rate of $10^{-4}$ is preferred to adapt to the 3D renderings split Vargas Hakim et al. (2023), as it represents a more aggressive shift.
Benchmarking.  We compare CLIPArTT with state-of-the-art methods. Specifically, we utilize adapted versions of TENT Wang et al. (2021) and LAME Boudiaf et al. (2022) tailored to CLIP. While the classifier logits are used in previous TTA research involving CNNs, we follow the standard practice and use the image-to-text similarity to obtain the final logits in CLIP. Only the visual encoder is optimized where needed. TENT now updates only the affine parameters in LN layers through entropy minimization directly on these logits. Notably, we employ 10 iterations for TENT adaptation, a choice informed by our findings, as evidenced in Fig 3. In LAME, the Laplacian regularizer is applied on the similarity between image features (obtained from the visual encoder). Finally, we include CLIP-tailored methods TTDN Zhou et al. (2023), TPT Shu et al. (2022), TDA Karmanov et al. (2024) and CALIP Guo et al. (2023). Both TDA and CALIP require hyperparameter tuning for optimal performance. To ensure fairness in our benchmarking process, we opted to fix these hyperparameters as recommended for the ImageNet dataset, as suggested in their respective works. The batch-size for TPT is reduced to 32 due to its reliance on image augmentations and high memory requirements.
@@
CIFAR-10 89.80 $\pm$0.05 90.04 $\pm$0.13 90.41 $\pm$0.07
CIFAR-10.1 85.37 $\pm$0.17 86.35 $\pm$0.27 86.07 $\pm$0.21
CIFAR-10-C 70.79 $\pm$0.15 71.17 $\pm$0.16 70.99 $\pm$0.15
@@
Iter = 1 Iter = 5 Iter = 10 Iter = 20
CIFAR-10 89.59 $\pm$0.01 90.54 $\pm$0.09 90.04 $\pm$0.13 88.32 $\pm$0.12
CIFAR-10.1 84.78 $\pm$0.02 86.67 $\pm$0.06 86.35 $\pm$0.27 84.33 $\pm$0.31
CIFAR-10-C 62.30 $\pm$0.06 68.75 $\pm$0.12 71.17 $\pm$0.16 70.55 $\pm$0.24
@@ 5.1 Ablation Studies
Number of classes for new prompts.  Determining the number of classes in building new prompts presents a critical aspect in our methodology. Table 1 illustrates various scenarios where $K\!=\!3$ emerges as a favorable choice, exhibiting a remarkable accuracy above $90\%$ across multiple instances of CIFAR-10-C datasets. This initial observation is further corroborated by the findings presented in Table 2. However, the decision becomes more nuanced when applied to CIFAR-100 datasets. While leveraging the top 3 classes contributes to enhanced accuracy, it does not guarantee optimal performance. Similarly, adopting a strategy akin to CIFAR-10, wherein $30\%$ of the classes are chosen, proves impractical due to resulting lengthy sentences that are impossible to tokenize. Nonetheless, our analysis shows that $K\!=\!3$ consistently yields superior performance, particularly evident in the average results across CIFAR-100-C. Consequently, we maintain $K\!=\!3$ as the preferred configuration for all subsequent experiments on CIFAR datasets.
Avg. Prompt Image Text Image + Text
CIFAR-10 76.94 $\pm$0.41 90.18 $\pm$0.02 89.05 $\pm$0.14 90.04 $\pm$0.13
CIFAR 10.1 67.83 $\pm$0.49 86.25 $\pm$0.37 84.85 $\pm$0.40 86.35 $\pm$0.27
CIFAR-10-C 43.08 $\pm$0.08 70.98 $\pm$0.15 70.03 $\pm$0.21 71.17 $\pm$0.16
CLIP BS = 16 BS = 32 BS = 64 BS = 128
CIFAR-10 88.74 85.89 $\pm$0.19 88.25 $\pm$0.15 89.48 $\pm$0.15 90.04 $\pm$0.13
CIFAR-10.1 83.25 81.55 $\pm$0.53 84.00 $\pm$0.31 85.40 $\pm$0.08 86.35 $\pm$0.27
CIFAR-10-C 59.22 64.72 $\pm$0.23 67.70 $\pm$0.23 69.82 $\pm$0.20 71.17 $\pm$0.16
CIFAR-100-C ImageNet-C
CLIP TENT TDA CLIPArTT CLIP TENT TDA CLIPArTT
Gaussian Noise 14.80 14.38 $\pm$0.14 08.20 $\pm$0.35 25.32 $\pm$0.14 12.27 12.28 $\pm$0.05 11.54 $\pm$0.05 20.18 $\pm$0.61
Shot noise 16.03 17.34 $\pm$0.27 09.58 $\pm$0.43 27.90 $\pm$0.05 12.20 08.46 $\pm$0.08 12.13 $\pm$0.13 20.94 $\pm$0.08
Impulse Noise 13.85 10.03 $\pm$0.13 07.63 $\pm$0.19 25.62 $\pm$0.09 12.89 15.71 $\pm$0.04 12.12 $\pm$0.08 19.95 $\pm$0.02
Defocus blur 36.74 49.05 $\pm$0.07 25.59 $\pm$0.41 49.88 $\pm$0.23 21.60 20.61 $\pm$7.05 21.39 $\pm$0.08 24.51 $\pm$0.13
Glass blur 14.19 03.71 $\pm$0.07 09.83 $\pm$0.56 27.89 $\pm$0.03 10.84 17.20 $\pm$0.08 10.74 $\pm$0.06 19.51 $\pm$0.02
Motion blur 36.14 46.62 $\pm$0.27 28.92 $\pm$0.18 47.93 $\pm$0.14 17.85 24.60 $\pm$0.04 18.45 $\pm$0.06 25.80 $\pm$1.16
Zoom blur 40.24 51.84 $\pm$0.15 31.08 $\pm$0.36 52.70 $\pm$0.06 16.38 21.13 $\pm$0.08 16.83 $\pm$0.03 24.00 $\pm$0.06
Snow 38.95 46.71 $\pm$0.21 32.94 $\pm$0.12 49.72 $\pm$0.01 21.91 24.19 $\pm$0.09 22.97 $\pm$0.05 27.26 $\pm$0.06
Frost 40.56 44.90 $\pm$0.27 34.84 $\pm$0.25 49.63 $\pm$0.12 23.33 23.34 $\pm$0.06 24.68 $\pm$0.06 26.56 $\pm$0.08
Fog 38.00 47.31 $\pm$0.04 31.13 $\pm$0.15 48.77 $\pm$0.04 26.39 28.58 $\pm$0.04 27.72 $\pm$0.05 34.68 $\pm$0.02
Brightness 48.18 60.58 $\pm$0.18 42.36 $\pm$0.10 61.27 $\pm$0.08 45.87 47.16 $\pm$0.01 48.06 $\pm$0.04 47.21 $\pm$0.07
Contrast 29.53 45.90 $\pm$0.11 18.03 $\pm$0.07 48.55 $\pm$0.24 16.16 21.62 $\pm$0.01 16.09 $\pm$0.05 23.38 $\pm$0.12
Elastic transform 26.33 33.09 $\pm$0.08 18.88 $\pm$0.24 37.45 $\pm$0.08 16.04 17.17 $\pm$0.01 17.75 $\pm$0.01 23.95 $\pm$0.04
Pixelate 21.98 26.47 $\pm$0.09 14.59 $\pm$0.30 33.88 $\pm$0.14 28.00 33.56 $\pm$1.71 30.03 $\pm$0.05 34.12 $\pm$0.09
JPEG compression 25.91 29.89 $\pm$0.07 17.56 $\pm$0.11 36.07 $\pm$0.32 27.44 31.89 $\pm$0.02 29.36 $\pm$0.01 32.23 $\pm$0.12
Average 29.43 35.19 22.08 41.51 20.61 23.17 21.32 26.95
Table 6: Accuracy (%) on CIFAR-100-C and ImageNet-C datasets with ViT-B/32 as visual encoder.
[…]
@@
Number of iterations.  In line with our approach for TENT, we investigate optimal number of iterations. As demonstrated in Table 3, iterating 10 times strikes the most favorable balance, exhibiting superior performance across the overall average on CIFAR-10-C and yielding optimal results for numerous corruption types. Nonetheless, it is imperative to acknowledge that employing a lower number of iterations (e.g., Iter = 5) may yield improved outcomes in scenarios with minimal or no distribution shift, while a higher number of iterations (e.g., Iter = 20) may be more effective in mitigating severe distribution shifts. As a result, 10 iterations are used for all forthcoming experiments.
@@
Performance of the model for different batch-sizes.  TTA methods have historically exhibited limitations when applied with small batch sizes. In our study, we investigate the performance of our model in this regard. As depicted in Table 5, performance improves significantly with increasing batch size. However, beyond a batch size of 8, no further performance gains are observed. This phenomenon can be attributed to the fact that, with smaller batch sizes, our method can potentially introduce uncertainty by using multiple classes, thereby diminishing performance, especially when the model is already confident. In subsequent experiments, we maintain a batch size of 128, consistent with prevailing practices in the state-of-the-art methodologies.
5.2 Comparison on different datasets
In standard TTA, the concept of domain shift is strictly related to the source dataset used for pretraining. In the context of CLIP, the presence of domain shifts is less evident, as the source data contained an enormous amount of images that likely include different domains. For this reason, we categorize our experiments per dataset type.
Shift CLIP LAME TENT TPT (BS=32) TTDN TDA CALIP CLIPArTT
CIFAR-10 88.74 89.36 $\pm$0.06 91.69 $\pm$0.10 88.06 $\pm$0.06 89.06 $\pm$0.02 84.09 $\pm$0.04 86.79  $\pm$0.01 90.04 $\pm$0.13
CIFAR-10.1 83.25 81.22 $\pm$0.33 87.60 $\pm$0.45 81.80 $\pm$0.27 83.77 $\pm$0.06 78.98 $\pm$0.37 81.02  $\pm$0.03 86.35 $\pm$0.27
CIFAR-100 61.68 58.27 $\pm$0.17 69.74 $\pm$0.16 63.78 $\pm$0.28 64.10 $\pm$0.05 60.32 $\pm$0.06 61.94 $\pm$0.01 69.79 $\pm$0.04
ImageNet 61.81 61.17 $\pm$0.02 62.13 $\pm$0.02 60.74 $\pm$0.16 61.85 $\pm$0.02 64.49 $\pm$0.01 58.50 $\pm$0.00 61.39 $\pm$0.09
VisDA-C (YT) 84.45 84.72 $\pm$0.01 84.41 $\pm$0.02 82.95 $\pm$0.01 83.29 $\pm$0.03 84.47 $\pm$0.08 85.02 $\pm$0.00 83.46 $\pm$0.03
PACS 93.65 93.68 $\pm$0.04 93.81 $\pm$0.03 93.23 $\pm$0.12 94.77 $\pm$0.07 92.94 $\pm$1.08 94.11 $\pm$0.00 93.95 $\pm$0.06
Office Home 77.53 75.46 $\pm$0.12 77.68 $\pm$0.06 77.20 $\pm$0.24 79.76 $\pm$0.04 77.70 $\pm$0.13 78.75 $\pm$0.01 77.56 $\pm$0.09
CIFAR-10-C 59.22 50.57 $\pm$0.30 67.56 $\pm$0.19 56.80 $\pm$0.22 60.13 $\pm$0.07 48.00 $\pm$0.68 56.58 $\pm$0.0 71.17 $\pm$0.16
CIFAR-100-C 29.43 26.23 $\pm$0.12 35.19 $\pm$0.16 30.46 $\pm$0.14 31.90 $\pm$0.04 22.08 $\pm$0.29 29.92 $\pm$0.01 41.51 $\pm$0.15
ImageNet-C 20.61 20.02 $\pm$0.16 23.17 $\pm$0.54 21.10 $\pm$0.06 21.28 $\pm$0.02 21.32 $\pm$0.06 21.00 $\pm$0.00 26.95 $\pm$0.12
VisDA-C (3D) 84.51 84.75 $\pm$0.17 83.85 $\pm$0.03 78.45 $\pm$0.05 80.45 $\pm$0.01 84.22 $\pm$0.10 83.81 $\pm$0.08 87.24 $\pm$0.04
Table 7: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with ViT-B/32 as visual encoder. The Shift column denotes a domain shift from standard images recognized by CLIP: corruption in CIFAR-10-C, CIFAR-100-C and ImageNet-C datasets, synthetic images in VisDA-C (3D).